OpenStudy (anonymous):

solve the equation 2(y-4)=4

OpenStudy (bradely):

Divide each side by 2 2(y-4)/2=4/2 y-4=2 Add 4 to each side y-4+4=2+4 y=6 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

2(y-4) = 4 First we can distribute that 2 to make it 2y - 8 = 4 Add 8 to both sides and we get 2y = 12 divide both sides by 2 and we get 7 = 6